# JWST spectroscopy of a blue binary cold classical Kuiper belt object

### v1

@@ 4 Summary
We presented JWST/NIRSpec observations of two binary cold classical KBOs. The measured reflectance spectra of the two targets differ significantly. 2001 XR254 belongs to the CH3OH-rich KBO compositional class, akin to other red cold classicals, with a steep continuum slope below 1.0 µm, strong signatures of aliphatic organics at 3.2–3.6 µm, deep features of CO2 and CO ices, and the distinctive presence of CH3OH-ice absorption bands between 2.2 and 2.6 µm. In contrast, the blue binary 2016 BP81 shows spectral features that largely coincide with the H2O-rich KBO compositional class. However, a detailed comparison of 2016 BP81 against the spectra of other H2O-rich KBOs revealed some notable differences, including a significantly stronger CO2-ice absorption feature at 4.27 µm, more prominent aliphatic organics signatures, and a higher band depth ratio between the 3- and 2-µm water-ice features. Along with the Neptune Trojan 2011 S0277 and the ETNO 2016 QV89, 2016 BP81 appears to delineate a distinct subclass within the broader H2O-rich KBO population, which may have originated in a separate formation environment from the rest of the H2O-rich KBO population.
@@
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #5940. The JWST data used in this paper can be found in MAST: http://dx.doi.org/10.17909/1wwx-1c65 (catalog 10.17909/1wwx-1c65).

### v2

@@ 4 Summary
We presented JWST/NIRSpec observations of two binary cold classical KBOs. The measured reflectance spectra of the two targets differ significantly: 2001 XR254 belongs to the organics-type KBO compositional class, akin to other red cold classicals, with a steep continuum slope below 1.0 m, strong signatures of aliphatic organics at 3.3–3.6 m, deep features of CO2 and CO ices, and the distinctive presence of CH3OH-ice absorption bands between 2.2 and 2.6 m; in contrast, the blue binary 2016 BP81 shows spectral features that largely coincide with the H2O-type KBO compositional class. However, a detailed comparison of 2016 BP81 against the spectra of other H2O-type KBOs revealed some notable differences, including a significantly stronger CO2-ice absorption feature at 4.27 m and a higher band-depth ratio between the 3 and 2 m H2O-ice features. Along with the Neptune Trojan 2011 S0277 and the ETNO 2016 QV89, 2016 BP81 appears to delineate a distinct cluster of objects with spectral properties that are broadly similar to those of a typical H2O-type KBO but nonetheless differ in several key ways that may be indicative of a separate formation environment from the rest of the H2O-type KBO population.
@@
Acknowledgments
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #5940. The JWST data used in this paper can be found in MAST: http://dx.doi.org/10.17909/1wwx-1c65 (catalog 10.17909/1wwx-1c65). , jwst (jwst), jwstspec (jwstspec), matplotlib (matplotlib), numpy (numpy), scipy (scipy).